- Problems during previous test 07Feb17:
- tx 3 and tx 6 had circuit breaker faults.
- on 13Feb 17 rf switches were installed at the broadband input
to switch between the control room synth value and a value input
locally at the transmitter.
What we did:
- mesh was at 5MHz position. used 5 MHz rcp , freq=5.125 MHz.
- 430 transmitter was used. data recorded with echotek and usrp.
- The interim correlator monitored the arcing using the ch
receiver. It also had the 430 tx sample in the spectra.
- 09:50 start to bring up tx at 5 MHz.
- 10:15 50KW each tx. stay here for about .5 hours
- sampled tx power much noisier than before . Jumps around by
- 10:54 80KW each tx. measured power still noisy .
- 11:12 tx 4 dropped out, then brought it back up.
- 11:31 found out about the rf switches being recently
- brought down tx2 and removed the rf switch.
- tx 2 power measurement now stable.
- 11:48 bring down tx1 . remove rf switch
- 11:45 started to bring up transmitters
- 11:55 bring down power all tx. remove the rest of the rf
- 12:07 all 6 tx back up to 80KW.
- leave the tx drive power fixed for 1.5 hours, see how the
- 12:24 tx2 jumps up by about .5 db.
- 12:48 some arcing seen on the ch monitor.
- turned off 430 transmitter. went under the bowl to check for
arcing.. nothing seen.
- 13:20 back from bowl. 430 tx turned back on.
- 14:03 tx 2 jumps down by 2+ db. then comes back a few
- 14:30 tx 2 jumps down by 3db then comes back up a few
- saw broadband go from 110 watts -> 85 watts -> 110
watts without changing the input drive levels.
- 14:40 adjust power levels in preparation for toggling...
about 75KW each tx.
- 14:47 toggling. 2 sec on, 4 sec off.
- 14:58 tx1 tripped with reverse power. reset.
- 15:07 back to cw. left power levels untouched to see
what the power levels were.
- they were a little higher then when we started..probably tx
cooler since 30% duty cycle.
- 15:10 tx2 dropped 80 to 45 KW. broadband went 110 to
80W. a few minutes later came back to 110 w
- 15:15 stopped.
Plots of output
and drive power:
The plots show the data we took
while the hf was on (.ps) (.pdf)
- The page has 3 frames:
- each color is a different transmitter (1 through 6).
- top frame: rf power (in dbm) measured at the output of
the rf coupler.
- middle frame: the same output power, but displayed in KWatts
- bottom frame:
- The synthesizer drive amplitude used
- The dashed green line shows when the rf was on
- the red line at the bottom is the total power from the
carriage house 430 receiver (polA 425-435 MHz). When we have
arcing, there are jumps in this power (rfi can also make
- 10.1 to 12.0 - you can see the jumping around of the output
- The rf switches were removed at 12.1.. things were a lot
more stable after that.
- 12.1 to 13.5 output drive power left constant.
- tx 2 red. is wandering around.
- 14.0,14.4 tx2 red drops output power level by up to 3db.
- problem was the broadband output was changing.
- 14.8 start auto toggle 2 on 4 off.
- 15.1 tx2 red drops output power level by 3db again..
comes back by itself.
- problem again with the broadband.
- 15.25 done
clp movie during hf test
The 430 clp data from the carriage was
processed and a movies were made. They showed how often
to the movies:
- ch at za=0, looking through the heated volume
- 10ms ipp, 440 usec rf pulse data taken for entire run.
- ch, 3 to 10 MHz were covered (upshifted)
- No data taken 12:47 to 13:23 (we were under the bowl looking
for arcs, so the 430 was turned off).
- ch, 25 mhz about 430 mhz.
- When processing the data
- the rfi removal algorithm was not used (didn't want to clip
the enhanced plasma line).
- Each frame of the movie was a 10 second integration.
- We saw lots of arcing in both the usrp and echotek
- 10 - 10:30 arcing when running at 50KW each tx
- no changes in drive level were made during this time.
- The morning also had the rf switches installed. After removing
the switches, the arcing continued.. so they probably did not
contribute to the arcing.
- 12 to 12:45 arcing got worse.. again no changes in drive level
- we then shut off the 430 tx to go under the dish and check
for arcing . nothing was found
- 13:24 430 back on. arcing continued.. but not quite as bad
- Images of the plasma line
- in both the usrp and the echotek data there is an image of
the plasma line shifted by +/- 5.125 MHz,. It is only there
when the heater is on.
- Since it is both usrp and echotek, we can blame it on a
particular backend. The harmonic/intermod must be occurring
upstairs (since we split the usrp,echotek at the crr2 output.
- tx4 tripped once, then came back up and ran reliably.
- tx2 output power jumps down by about 3db for a few minutes and
then comes back to previous value.
- saw the output of the broadband change each time.
- This happened 3 times during the test.
- saw some arcing on the ch monitor. On the scope these looked
like shorter pulses. So the interim correlator 1 sec averages
didn't see them.
- We see arcing in the clp movies
- There is a harmonic/intermod of the plasma line offset
by +/- 5.125 MHz. This is seen in the usrp and the
<- page up